OpenStudy (kamille):

1.5 microM = 1,5*10^-3mM is this correct?

OpenStudy (kamille):

@ash2326

OpenStudy (kamille):

and can you check this one too: \[5nmol=5*10^{-9}p mol= ... mol?\]

OpenStudy (phi):

1.5 microM = 1,5*10^-3mM is this correct? yes \[ 1.5 \ \mu M = 1.5\cdot 10^{-3} \ mM\] where mM means milli M (whatever M stands for)

OpenStudy (kamille):

yes, m means mili, can you check the one with p(pico) and how should I convert just to normal moles/?

OpenStudy (phi):

see http://www.telecomabc.com/p/prefix.html

OpenStudy (kamille):

I know what it means, I just don't know math (how to convert from one to another)

OpenStudy (phi):

5 n mol = 5 * 10^-9 moles 1 p mol = 1 * 10^-12 moles in other words a p mol is 1000 times smaller than a n mol or n mol is 1000 times bigger than a p mol

OpenStudy (kamille):

Oh, okay, now I get this. Thanks!

OpenStudy (phi):

5 n mol = 5000 p mol